# Embodied World Models for Decision Making

### v3

@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled sophisticated multi-agent systems for handling task plans in embodied AI. Embodied agents deployed in real-world environments are often tasked with executing extended sequences of actions in pursuit of high-level goals. For such agents to be effective, their plans must be accurate in terms of task completion, efficient in their use of time and actions, and safe in avoiding undesirable or unintended consequences. With the advent of LLMs, there has been growing interest in using these systems to generate and reason about structured plans from natural language input. Recent work has demonstrated that LLMs can decompose high-level goals into executable sub-goals and, in multi-agent setups, evaluate and revise each other’s outputs through structured dialogue ([1],  [2]).
Despite this progress, LLM-generated plans have been shown to be not perfect [3, among others], and need to be verified. Furthermore, the quality of the datasets used to train planning for embodied agents remains a significant bottleneck. Many widely used corpora, such as TEACh ([4]), a dataset of household interactions collected via human operation, contain a large number of suboptimal behaviors ([5]). These include, for example, picking up irrelevant or unnecessary items, leaving goals unaccomplished due to erroneous instruction, or toggling appliances that were never activated. While such issues result in unnecessarily lengthy plans and introduce noise in learning signals during imitation or reinforcement learning, they also present an underexplored opportunity: human demonstrations naturally contain error-recovery sequences that showcase how humans recognize and correct mistakes in real-time. Our verification framework, by explicitly identifying these erroneous actions, not only enables cleaner training data through removal but also facilitates future research into learning from human error recovery patterns, potentially yielding agents that exhibit human-like resilience and adaptability when facing unexpected situations or their own execution errors.
@@ 2.1 LLMs as Plan Generators
In the earlier work, Gella et al. have demonstrated the dialogue act prediction based approach using the TEACh dataset [6]. This was a classification based approach trying to predict the next dialogue act given the action history and environment. They have expanded this work so as to use an LLM based planning approach in [3] to generate an entire action sequence from a high-level goal. Another work by Huang et al. [7] demonstrated that while large pre-trained models contain sufficient world knowledge to produce semantically plausible plans, these plans are often not executable because they fail to map onto an agent’s specific, admissible actions. This fundamental gap between plausible text and executable actions highlights the core challenge of grounding.
Subsequent research has focused on closing this gap by providing LLMs with stronger environmental context. For instance, ProgPrompt [8] structures the prompt as a Python program, providing the LLM with available actions as import statements and objects as a list, thereby constraining the generated output to be more syntactically correct and environmentally aware. Similarly, LLM-Planner [9] couples few-shot prompting with online simulator feedback, appending error traces to the prompt to enable replanning when a step fails. Other methods, like TaPA [10], inject explicit world state, such as an inventory table, directly into the prompt to reduce object hallucination. Our work is orthogonal to these plan generation methods; instead of improving the initial output, we focus on creating a robust post-hoc verification process to clean noisy, human-authored trajectories, ensuring that any downstream planner inherits a leaner and less ambiguous set of demonstrations.
@@ 2.2 Verification and Refinement
Given the difficulties in direct plan generation, a powerful alternative paradigm has emerged: using LLMs to critique, verify, and refine plans within an iterative loop. This approach often decomposes the problem into distinct agentic roles. The Reflexion framework, for example, splits an agent into an Actor, Evaluator, and Self-Reflection module, using verbal reinforcement to iteratively improve its policy without gradient updates [1]. Our Judge-Planner framework externalizes this critique process into an independent, two-agent system, which facilitates explicit precision-recall accounting and allows for modular substitution of the Judge LLM. This multi-agent perspective is echoed in broader frameworks like AutoGen, which coordinates specialized conversable agents [2], and CoELA, which pairs a "Commander" LLM with domain-specific modules for robot teams [11].
More directly, recent work has built systems specifically for pre-execution plan verification. VerifyLLM [12] proposes a verification architecture that analyzes action sequences to identify logical inconsistencies, missing prerequisites, and redundancies, using Linear Temporal Logic (LTL) as a formal intermediate representation to guide the LLM’s analysis. The principle of using an LLM to verify an intermediate output is not limited to robotics; LLatrieval [13] employs an LLM to iteratively verify and refine the documents retrieved by a RAG system, demonstrating the general power of this verification-and-update pattern. Our framework contributes to this line of work by demonstrating that effective, high-recall verification can be achieved using zero-shot natural language critique alone, without requiring formal methods like LTL.
@@
Casting an LLM in a verification role makes it an instance of an "LLM-as-a-Judge," a rapidly growing field with its own set of tools and challenges. A systematic analysis by Li et al. [14] formally evaluates the distinct contributions of LLMs as solvers, verifiers, and heuristic functions, concluding that LLMs are often more effective at providing comparative feedback than at generating correct solutions from scratch – a finding that strongly motivates our verification-centric approach.
However, the reliability of LLM judges is not guaranteed. Research into their adversarial robustness has shown that simple, universal triggers can inflate judge scores, highlighting the fragility of naive evaluation [15]. To promote standardized evaluation, public benchmarks such as MT-Bench and the Chatbot Arena have been developed to provide open, crowd-validated tests where models like GPT-4 have achieved human-level agreement [16]. Our work adopts their strict evaluation protocols. Furthermore, systematic surveys have begun to chart the field, offering taxonomies of judgment tasks and techniques for bias mitigation [17, 18], which inform the design of our evaluation suite. Finally, within the embodied domain, benchmarks like PARTNR provide large-scale, multi-agent tasks with built-in verification checkpoints [19]. We instead target the TEACh dataset for its organic, noisy human demonstrations, which better approximate the data our framework is designed to clean.
@@ 5 Conclusion
These results establish plan verification as a distinct, language-level capability of LLMs, with strong implications for both dataset curation and agent performance in downstream learning tasks. Future work will extend this framework to vision-grounded settings, explore adversarial robustness of judge prompts, and test the framework’s scalability on larger multi-agent corpora such as PARTNR [19].

### v4

@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled sophisticated multi-agent systems for handling task plans in embodied AI. Embodied agents deployed in real-world environments are often tasked with executing extended sequences of actions in pursuit of high-level goals. For such agents to be effective, their plans must be accurate in terms of task completion, efficient in their use of time and actions, and safe in avoiding undesirable or unintended consequences. With the advent of LLMs, there has been growing interest in using these systems to generate and reason about structured plans from natural language input. Recent work has demonstrated that LLMs can decompose high-level goals into executable sub-goals and, in multi-agent setups, evaluate and revise each other’s outputs through structured dialogue ([shinn2023reflexion],  [wu2023autogen]).
Despite this progress, LLM-generated plans have been shown to be not perfect [padmakumar-etal-2023-multimodal, among others], and need to be verified. Furthermore, the quality of the datasets used to train planning for embodied agents remains a significant bottleneck. Many widely used corpora, such as TEACh ([padmakumar2021teach]), a dataset of household interactions collected via human operation, contain a large number of suboptimal behaviors ([min2022dontcopyteacher]). These include, for example, picking up irrelevant or unnecessary items, leaving goals unaccomplished due to erroneous instruction, or toggling appliances that were never activated. While such issues result in unnecessarily lengthy plans and introduce noise in learning signals during imitation or reinforcement learning, they also present an underexplored opportunity: human demonstrations naturally contain error-recovery sequences that showcase how humans recognize and correct mistakes in real-time. Our verification framework, by explicitly identifying these erroneous actions, not only enables cleaner training data through removal but also facilitates future research into learning from human error recovery patterns, potentially yielding agents that exhibit human-like resilience and adaptability when facing unexpected situations or their own execution errors.
@@ 2.1 LLMs as Plan Generators
In the earlier work, Gella et al. have demonstrated the dialogue act prediction based approach using the TEACh dataset [gella2022dialog]. This was a classification based approach trying to predict the next dialogue act given the action history and environment. They have expanded this work so as to use an LLM based planning approach in [padmakumar-etal-2023-multimodal] to generate an entire action sequence from a high-level goal. Another work by huang2022language demonstrated that while large pre-trained models contain sufficient world knowledge to produce semantically plausible plans, these plans are often not executable because they fail to map onto an agent’s specific, admissible actions. This fundamental gap between plausible text and executable actions highlights the core challenge of grounding.
Subsequent research has focused on closing this gap by providing LLMs with stronger environmental context. For instance, ProgPrompt [singh2022progpromptgeneratingsituatedrobot] structures the prompt as a Python program, providing the LLM with available actions as import statements and objects as a list, thereby constraining the generated output to be more syntactically correct and environmentally aware. Similarly, LLM-Planner [song2023llmplanner] couples few-shot prompting with online simulator feedback, appending error traces to the prompt to enable replanning when a step fails. Other methods, like TaPA [wu2023tapa], inject explicit world state, such as an inventory table, directly into the prompt to reduce object hallucination. Our work is orthogonal to these plan generation methods; instead of improving the initial output, we focus on creating a robust post-hoc verification process to clean noisy, human-authored trajectories, ensuring that any downstream planner inherits a leaner and less ambiguous set of demonstrations.
@@ 2.2 Verification and Refinement
Given the difficulties in direct plan generation, a powerful alternative paradigm has emerged: using LLMs to critique, verify, and refine plans within an iterative loop. This approach often decomposes the problem into distinct agentic roles. The Reflexion framework, for example, splits an agent into an Actor, Evaluator, and Self-Reflection module, using verbal reinforcement to iteratively improve its policy without gradient updates [shinn2023reflexion]. Our Judge-Planner framework externalizes this critique process into an independent, two-agent system, which facilitates explicit precision-recall accounting and allows for modular substitution of the Judge LLM. This multi-agent perspective is echoed in broader frameworks like AutoGen, which coordinates specialized conversable agents [wu2023autogen], and CoELA, which pairs a "Commander" LLM with domain-specific modules for robot teams [zhang2023coela].
More directly, recent work has built systems specifically for pre-execution plan verification. VerifyLLM [grigorev2025verifyllm] proposes a verification architecture that analyzes action sequences to identify logical inconsistencies, missing prerequisites, and redundancies, using Linear Temporal Logic (LTL) as a formal intermediate representation to guide the LLM’s analysis. The principle of using an LLM to verify an intermediate output is not limited to robotics; LLatrieval [li2024llatrieval] employs an LLM to iteratively verify and refine the documents retrieved by a RAG system, demonstrating the general power of this verification-and-update pattern. Our framework contributes to this line of work by demonstrating that effective, high-recall verification can be achieved using zero-shot natural language critique alone, without requiring formal methods like LTL.
@@
Casting an LLM in a verification role makes it an instance of an "LLM-as-a-Judge," a rapidly growing field with its own set of tools and challenges. A systematic analysis by li2024systematic formally evaluates the distinct contributions of LLMs as solvers, verifiers, and heuristic functions, concluding that LLMs are often more effective at providing comparative feedback than at generating correct solutions from scratch – a finding that strongly motivates our verification-centric approach.
However, the reliability of LLM judges is not guaranteed. Research into their adversarial robustness has shown that simple, universal triggers can inflate judge scores, highlighting the fragility of naive evaluation [raina2024robust]. To promote standardized evaluation, public benchmarks such as MT-Bench and the Chatbot Arena have been developed to provide open, crowd-validated tests where models like GPT-4 have achieved human-level agreement [zheng2023mtbench]. Our work adopts their strict evaluation protocols. Furthermore, systematic surveys have begun to chart the field, offering taxonomies of judgment tasks and techniques for bias mitigation [gu2025survey, li2025generation], which inform the design of our evaluation suite. Finally, within the embodied domain, benchmarks like PARTNR provide large-scale, multi-agent tasks with built-in verification checkpoints [chang2022partnr]. We instead target the TEACh dataset for its organic, noisy human demonstrations, which better approximate the data our framework is designed to clean.
@@ 5 Conclusion
These results establish plan verification as a distinct, language-level capability of LLMs, with strong implications for both dataset curation and agent performance in downstream learning tasks. Future work will extend this framework to vision-grounded settings, explore adversarial robustness of judge prompts, and test the framework’s scalability on larger multi-agent corpora such as PARTNR [chang2022partnr].